Blue Ridge Marketing Inc. manufactures two products, A and B. Presently, the company uses a single plantwide factory overhead rate for allocating overhead to products. However, management is considering moving to a multiple department rate system for allocating overhead. The following table presents information about estimated overhead and direct labor hours.


Overhead
Direct
Labor Hours (dlh)
Product
A B
Painting Dept. $241,900 11,500 dlh 2 dlh 12 dlh
Finishing Dept. 72,100 11,300 5 7
    Totals $314,000 22,800 dlh 7 dlh 19 dlh

Using a single plantwide rate, determine the overhead rate per unit for Blue Ridge Marketing Inc.'s Product B.

$42.07

$96.39

$261.63

$13.77

The Ramapo Company produces two products, Blinks and Dinks. They are manufactured in two departments, Fabrication and Assembly. Data for the products and departments are listed below.


Product
Number of
Units
Labor Hours
Per Unit
Machine Hours
Per Unit
Blinks 957 2 4
Dinks 1,986 3 6

2)All of the machine hours take place in the Fabrication department, which has an estimated overhead of $108,200. All of the labor hours take place in the Assembly department, which has an estimated total overhead of $79,800.

The Ramapo Company uses a single overhead rate to apply all overhead costs. What would the single plantwide rate be if it was based on machine hours instead of labor hours?

a.$27.46 per machine hour

b.$9.55 per machine hour

c.$14.93 per machine hour

d.$11.94 per machine hour
Answer #1

1) Overhead rate = Estimated overhead/estimated labor hour = 314000/22800 = 13.77 per labor hour

Overhead applied to product B = 13.77*19 = 261.63

So answer is c) $261.63

2) Total machine hour = (957*4+1986*6) = 15744 Machine hour

Overhead rate = (108200+79800)/15744 = 11.94

So answer is d) $11.94 per machine hour